# Learning to Explore and Select for Coverage-Conditioned Retrieval-Augmented Generation

### v1

@@ 3.4.1 Parsing Outlines
In this stage, LLM sequentially extracts candidate outlines from QTree that satisfies instructions of C2 queries [^3] Our preliminary verification identifies that sequentially generating candidate outlines shows more diversity than temperature-based sampling. Refer to Appendix B for case studies.. Table 3 illustrates the example of the candidate outline, providing a hierarchical subset of queries (i.e., plot - main event - conflict & climax) about The Woman Hunt. We set the number of candidate outlines to three per each C2 query.
@@
We fix the number of queries within each outline to four, guaranteeing that all queries are directly connected or neighboring within QTree, as illustrated in Table 3. Additionally, we verify the JSON parsability of each outline and ensure that all queries do not overlap each other. For the efficient usage of API calls, we heuristically remove queries in leaf nodes if an outline contains more than four queries and include the outline as a candidate.
@@ 3.4.2 Evaluating Outline Quality
In order to evaluate the quality of candidate outlines, we leverage LLM (gpt-4-0125-preview) to serve as a judge deciding whether the content on each candidate outline follows C2 query. More precisely, we prompt the model to assign scores, measuring how faithfully the outline aligns with the C2 query. The scoring mechanism employs a five-point Likert scale, along with a corresponding rationale generation for the evaluation. Since outlines have significantly shorter context than long-form text while containing core contents (Kellogg, 1988), it is expected that the evaluation of the LLM shows better alignment with human intuition than directly evaluating long context.
@@ 3.5 Training QPlanner (Step 3)
To deal with arbitrary C2 queries and generate corresponding outlines, we train a language model named QPlanner. We instruct QPlanner to sequentially generate QTree and selected outline, as we intend that QTree serves like an intermediate Chain-of-Thought (Wei et al., 2022b) reasoning process.
In the training phase, we employ supervised fine-tuning (SFT) and alignment tuning for the training QPlanner. First, we train the Llama-2-7B-Chat model (Touvron et al., 2023) using 10K C2 queries mapped with 31K candidate outline pairs, which are constructed through Section 3.4.1. This training phase allows the model to generate formatted outlines following C2 queries. Then we further align the preferred outline by adopting a variant of direct preference optimization (DPO) (Rafailov et al., 2023). Following Tunstall et al. (2023) that show the possibility of distilling the preference of large teacher models into the targeted model, we utilize LLM evaluation previously acquired in Section 3.4.2 as a reward signal for aligning QPlanner. We regard the chosen sample as the highest-scored and the rejected sample as the lowest-scored outline. We do not use samples whose highest and lowest scores are the same in the alignment stage. The amount of the final training sample is 31,488 in SFT and 8,568 in DPO, respectively. Refer to Appendix F for further details, such as hyperparameters.
@@ 4 Experiments
Table 4: Mean and standard deviation (SD) for automatic outline evaluation (five-point Likert scale). Our DPO-QPlanner scores the highest mean score and the lowest SD, indicating robust improvement.
4.1 Automatic Evaluation
We let LLM (gpt-4-0125-preview) conduct automatic scoring for outlines on our test set. For long-form response evaluation, we do not automatically measure due to the lack of reliability in long-form text evaluation (Xu et al., 2023).
@@ 4.1.1 Mean Score Comparison
Table 4 shows our test result with a five-point Likert scale. We score outlines generated by each trained model [^4] A few cases return outline with 3 or 5 queries, which is not an ideal number of the output (i.e., 4), but we do not filter them in our evaluation., focusing on whether the content of outlines follows given C2 queries (as mentioned in Section 3.4.2).
DPO-QPlanner (Ours)
While our SFT model shows a higher score than the random baseline, we find that preference alignment based on the SFT model using our candidate outlines, which are ranked in Section 3.5, significantly improves the evaluation score. It should be noted that our reward signal is based on the language model’s generation whose subjectivity is similar to human’s, different from Ma et al. (2023); Yoon et al. (2024) whose reward signals are more likely to be an explicit and gold criterion (  e.g., retrieved ranking of the target document). Since there is no proper baseline to verify the effectiveness of our outlining framework, we construct the following baselines as an initial point to estimate performance.
Random Baseline
Since the output of QPlanner includes QTree, we can extract arbitrary outline regardless of the C2 queries. Specifically, we select a random background query from QTree generated by our SFT-QPlanner, then extend the branch to randomized directions (i.e., upper depth, neighbor, or lower depth) until four queries are connected as a single outline [^5] Intent operation is also not considered in the random baseline.. We find that the random baseline shows the lowest mean score (2.57) with the highest standard deviation (1.44) on our test set.
DPO-SynNeg
To further explore the effectiveness of selected negative samples in DPO-QPlanner, we prepare another DPO model trained with different types of negative samples. While negative samples of DPO-QPlanner are generated by LLM, as described in Section 3.5, it is also possible to heuristically synthesize negative samples using QTree, background query, and intent operation. This procedure is similar to generating random baseline, except for (1) setting the starting point as the background query and (2) ensuring synthesized outlines have the opposite intent operation to the original intent. That is, the intent of synthetic negative (SynNeg) outlines explicitly conflicts with that of the positive (chosen) outlines, while the intent of negative outlines of DPO-QPlanner are aligned with that of positive outlines. On the 8K DPO-QPlanner training set, we maintain the chosen (i.e., highest-scored) outlines and replace rejected outlines with synthetically generated negative outlines.
@@
On the comparison with DPO-QPlanner and DPO-SynNeg, we find that negative samples of DPO-QPlanner are notably effective than the other one. We conjecture that constructing “hard” negatives [^6] We define LLM-generated negative samples as “hard” when compared to synthetic negative samples whose intent operations are explicitly opposite to positive outlines. (i.e., less scored subtrees with the “same” intent) is an important factor to align with C2 queries, as discussed in previous studies (Rosset et al., 2023; Scarlatos et al., 2024).
DPO-Combined
We also measure the performance of combining negative samples in DPO-QPlanner and DPO-SynNeg. While the test score slightly increases compared to DPO-SynNeg, performance becomes worse than DPO-QPlanner despite the doubled amount of training samples.
4.1.2 Pairwise Score Comparison
We also measure pairwise comparison between models in Figure 3. Our DPO-QPlanner gets the highest score on the same C2 query, which is intuitively aligned with the result in Table 4. Given the fact that exhaustively devising preferred outline for C2 queries is labor-intensive, our QPlanner is an effective solution for exploring and providing preferred outlines for composing long-form responses.
@@ 4.1.2 Pairwise Score Comparison
Figure 3: Pairwise score comparison per each C2 query in automatic outline evaluation. Outlines of our DPO-QPlanner are judged as the best among all other baselines.
4.2 Human Evaluation
@@
We also conduct a human study to identify the effectiveness of QPlanner. We describe detailed information in Appendix G such as demographics and payment.
4.2.1 Outline Evaluation
@@
Recruited evaluators read and rate two outlines (SFT- and DPO-QPlanner) for randomly selected 100 C2 queries from the test set using the same five-point Likert scale criteria as the automatic evaluation. Each evaluator rates from 4 to 10 outlines in random order, and each outline has at least 6 evaluators (Average: 6.55, Max: 15). We intend to have as many evaluators as possible to rate individual outline to gather a collective rating for each set. This is because even with simple outlines, judging if outlines with unfamiliar subjects successfully follow C2 queries is highly intellectual and unavoidably subjective task.
@@
As expected for this type of subjective task, evaluators do not reach a high inter-rater reliability (Krippendorff’s $\alpha$=0.22 for SFT ratings 0.23 for DPO-QPlanner ratings between human raters). However, more importantly, we find significant positive correlations between human-rated scores and LLM-rated scores – both for SFT (Pearson’s $r$ = 0.51, p-value $<$ 0.001) and DPO-QPlanner (Pearson’s $r$ = 0.39, p-value $<$ 0.001), which indicates positive relationships with large and medium strength, respectively. Moreover, DPO-QPlanner receives higher scores (Mean= 3.29, Std=0.81) than SFT (Mean= 3.03, Std=0.78). This trend is supported even when we regress scores on model version (SFT or DPO-QPlanner) and the total length of outlines in characters, with outline id as a fixed effect (Model:$b$=0.27, p-value=0.01; Outline length: $b$=0.01, p-value=0.30). That is, the length of outlines is not predictive of scores.
4.2.2 Response Evaluation
@@
In addition to outline evaluation, we recruit another participants to evaluate 100 randomly sampled RAG responses. Ten evaluators are assigned for each response. According to Kim et al. (2023), the writing format of model responses affects human preferences. Therefore, we fix the format with Markdown to compare responses by focusing only on their content. In addition, we prompt LLM to generate responses by strictly relying on given evidence to prevent LLM from arbitrarily responding with its parametric knowledge. We provide a description of the document retrieval procedure in Section G.2.1.
@@
Regarding evaluation criteria, we first instruct participants to judge whether generated responses follow requirements within C2 queries or not (Query Satisfaction in Figure 4). We guide to annotate "Yes" if responses at least partially handle topics within C2 queries. For response pairs annotated as "Yes" in both models, participants select their preferred response from the two models (Response Preference in Figure 4).
@@
We verify whether queries within outlines can help search relevant documents. We compare responses of vanilla RAG with those of QPlanner-DPO using the exactly same prompt. In other words, outline queries of QPlanner-DPO only affect search results. As illustrated in Figure 4(a), we find that conventional RAG pipeline does not properly handle C2 queries, whose requirements are far more complicated than normal queries. Furthermore, among responses that satisfy the requirements of C2 queries, responses of QPlanner-DPO are mostly preferred.
@@
We further explore whether better outlines lead to better responses. In this setup, we compare responses of QPlanner-SFT with those of QPlanner-DPO. Likewise, the exactly same prompt is used for the comparison and queries within outlines are included to compose responses, along with document searching. Results in Figure 4(b) indicate that QPlanner aligned with preference outlines tends to provide preferred outlines, although QPlanner-SFT also shows a similar tendency with QPlanner-DPO, which is different from that of Vanilla-RAG.
@@
Figure 4: Human evaluation results
For both studies, we conduct two McNemar’s tests (McNemar, 1947) to examine whether the differences we found are statistically significant. We choose to run McNemar’s tests because they are suited for our experiment settings, where the outcome is binary (i.e., satisfactory or not), and each human evaluator judges both responses for the same query (within-subjects design). The contingency tables used for the tests can be found in Table 10. The results confirm that QPlanner-DPO outperforms Vanilla RAG (test-statistics=$60$, p-value<0.001), and QPlanner-DPO outperforms QPlanner-SFT (test-statistics=$207$, p-value=0.004).
@@ 5 Conclusion
Customizing information coverage within C2 queries requires complex selection of subtopics to compose long-form responses. We investigate the role of query outline in terms of assistive search queries and intermediate content draft for RAG task. To simulate C2 scenarios, we construct QTree, hierarchical sets of queries representing diverse perspectives of the original query. Subsequently, we train QPlanner which extracts customized outlines from QTree for C2 queries. Our findings based on automatic and human evaluation show that (1) preference-aligned QPlanner can provide better outlines, (2) outlines enable searching more relevant documents, and (3) well-generated outlines lead to preferred responses. We believe our work shows the possibility of QTree as a testbed for exploring effective pre-writing strategy to deal with complicated queries.
Limitation
We fix the number of queries in each outline to four, but it can be adjusted (i.e., composing more or less queries) according to the topic. We plan to adapt outlines in a flexible manner in our future work.
It should be noted that retrieved documents in our RAG setup can also affect final responses. Although we set the same searching configuration among all methodologies, additional fact verification is still needed for practical applications.
In addition, despite demonstrating significant performance gain both in automatic and human evaluation, we find that state-of-the-art LLM still has difficulty in generating long-form responses that handle detailed coverage of C2 query. This is presumably due to the complexness of C2 queries, but it arises the importance of constructing meticulous benchmark evaluating long-form responses of complicated queries. This will be another direction of our future work.

### v2

@@ 3.4.1 Parsing Outlines
In this stage, LLM sequentially extracts JSON-formatted candidate outlines from QTree that satisfy instructions of C2 queries [^4] Our preliminary verification identifies that sequentially generating candidate outlines shows more diversity than temperature-based sampling. Refer to Appendix B for case studies.. Table 3 visualizes the example of a candidate outline, consisting of hierarchical subqueries (i.e., plot - main event - conflict & climax) about The Woman Hunt. We extract three candidate outlines per each C2 query.
@@
We fix the number of subqueries within each outline to four, guaranteeing that all subqueries are directly connected or neighboring within QTree, as illustrated in Table 3. Additionally, we verify the JSON parsability of each outline and ensure that all subqueries do not overlap each other. For the efficient usage of API calls, we heuristically remove subqueries in leaf nodes if an outline contains more than four subqueries and include the outline as a candidate.
@@ 3.4.2 Evaluating Outline Quality
In order to rank three candidate outlines, we leverage LLM (gpt-4-0125-preview) to serve as a judge deciding whether the content on each candidate outline follows C2 query. More precisely, we prompt the model to assign five-point Likert-scale scores with rationales, measuring how faithfully the outline aligns with the C2 query. Since outlines are significantly shorter than long-form text while maintaining core contents (Kellogg, 1988), it is expected that evaluating outlines is more efficient and intuitive than directly evaluating long responses. These scored outlines are utilized as supervision and alignment pairs for training QPlanner, which will be described in further sessions.
@@ 3.5 Training QPlanner (Step 3)
To generalize with arbitrary C2 queries, we train a 7B language model named QPlanner. We instruct QPlanner to sequentially generate QTree and select an outline, as we intend that QTree serves like an intermediate Chain-of-Thought (Wei et al., 2022) reasoning process. More technical details are described in Section 4.1.
@@ 4 Experiments
Table 4: Mean and standard deviation (SD) for automatic outline evaluation (five-point Likert scale). DPO-QPlanner scores the highest mean score and the lowest SD, indicating robust improvement.
4.1 Training Details of QPlanner
We employ supervised fine-tuning (SFT) and alignment tuning for the training QPlanner. First, we train the Llama-2-7B-Chat model (Touvron et al., 2023) using 10K C2 queries mapped with 31K candidate outline pairs, constructed through Section 3.4.1. This training phase allows the model to generate formatted outlines following C2 queries (named SFT-QPlanner hereafter).
Then we further align the preferred outline by adopting a variant of direct preference optimization (DPO) (Rafailov et al., 2023). Following Tunstall et al. (2024) that show the possibility of distilling the preference of large teacher models into a targeted model, we utilize LLM evaluation scores previously acquired in Section 3.4.2 as reward signals for aligning QPlanner (named DPO-QPlanner hereafter). We regard the highest-scored outline as a positive (chosen) sample and the lowest-scored outline as a negative (rejected) sample. We skip samples whose highest and lowest scores are the same in the alignment stage.
The amount of the final training sample is 31,488 for SFT-QPlanner and 8,568 for DPO-QPlanner, respectively. Refer to Appendix F for further details, such as hyperparameters.
4.2 Baselines for Outline Comparison
Random Baseline
Since the output of QPlanner accompanies QTree as an intermediate reasoning process, we can extract an arbitrary outline by leveraging this, regardless of the C2 queries. Specifically, we select a random background subquery from QTree generated by SFT-QPlanner, then extend the branch to randomized directions (i.e., upper depth, neighbor, or lower depth) until four subqueries are connected as a single outline. Intent operation is not considered in this random baseline.
DPO-SynNeg
To further explore the effectiveness of selected (i.e., LLM-scored) negative samples in DPO-QPlanner, we prepare another DPO model trained with different types of negative samples. While negative samples of DPO-QPlanner are based on LLM scores, we can also heuristically synthesize negative samples with QTree, background subquery, and intent operation. This procedure is similar to generating random baseline, except for ensuring that synthesized outlines have the opposite intent operation to the original intent. For example, if the positive outline includes background subquery, the synthesized outline is designed to exclude that subquery by selecting another random background subquery within QTree. For the opposite situation, the synthesized outline must contain background subquery of the positive outline. On the 8K DPO-QPlanner training set, we maintain the positive samples and replace negative samples with synthetically generated outlines.
DPO-Combined
We also measure the performance of combining negative samples in DPO-QPlanner and DPO-SynNeg. That is, the number of training samples is doubled.
5 Results
5.1 Automatic Outline Evaluation
We prompt LLM (gpt-4-0125-preview) to score generated outlines in the test set. We use scoring rubric in Table 9 in Section G.1.
@@ 5.1.1 Mean Score Comparison
Table 4 shows our test result with a five-point Likert scale. We score outlines generated by each trained model [^5] A few cases return an outline with 3 or 5 queries, which is not an ideal number of the output (i.e., 4), but we do not filter them in our evaluation., focusing on whether the content of outlines follows given C2 queries (as mentioned in Section 3.4.2).
We find that the random baseline shows the lowest mean score (2.57) with the highest standard deviation (1.44) on our test set. While SFT-QPlanner shows a higher score than the random baseline, we find that DPO-QPlanner significantly improves the score (3.16) with the lowest standard deviation (1.33). It implies that we can leverage LLM-generated scores as reward signals in query outlining when constructing positive-negative pairs, even in the absence of explicit and gold reward criteria for their construction (Ma et al., 2023; Yoon et al., 2024).
@@
On the comparison with DPO-QPlanner and DPO-SynNeg, we observe that negative samples of DPO-QPlanner are notably more effective than the other one, since the only difference between them is the type of negative samples. We conjecture that constructing “hard” negatives [^6] We define LLM-generated negative samples as “hard” when compared to synthetic negative samples whose intent operations are explicitly opposite to positive outlines. (i.e., less scored subtrees with the “same” intent) is an important factor to align with C2 queries, sharing insights with different studies on hard negatives (Rosset et al., 2023; Scarlatos et al., 2024). Regarding DPO-Combined, the performance becomes worse than DPO-QPlanner despite the doubled amount of alignment pairs, implying the importance of selective negative samples.
@@
Figure 3: Pairwise comparison for each C2 query in automatic outline evaluation.
5.1.2 Pairwise Comparison
For comprehensive evaluation, we also compare pairwise scores among models. As illustrated in Figure 3, generated outlines of DPO-QPlanner are more preferred than all other baselines on the same C2 query, which is aligned with the atomic scoring result in Table 4. Given the fact that exhaustively devising preferred outlines for C2 queries is labor-intensive, our QPlanner is an effective solution for exploring and creating preferred outlines for long-form responses.
5.2 Human Outline Evaluation
@@
We conduct a human study to identify the effectiveness of QPlanner. We describe detailed experimental setup in Appendix G, such as demographics and payment.
@@
We let participants read and rate two outlines generated from SFT- and DPO-QPlanner for randomly selected 100 C2 queries from the test set using the same five-point Likert scale criteria as the automatic evaluation. Each evaluator rates from 4 to 10 outlines in random order (to avoid position bias), and each outline has at least 6 evaluators (Average: 6.55, Max: 15). We intend to have as many evaluators as possible to rate individual outlines to gather a collective rating for each set. This is because even with simple outlines, judging outlines with unfamiliar topics is a highly intellectual and unavoidably subjective task.
@@
Consequently, we find significant positive correlations between human-rated scores and LLM-rated scores – both for SFT (Pearson’s $r$ = 0.51, p-value $<$ 0.001) and DPO-QPlanner (Pearson’s $r$ = 0.39, p-value $<$ 0.001), which indicates positive relationships with large and medium strength, respectively. Moreover, DPO-QPlanner receives higher human scores (Mean= 3.29, Std=0.81) than SFT (Mean= 3.03, Std=0.78) [^7] This trend is supported even when we regress scores on model version (SFT or DPO-QPlanner) and the total length of outlines in characters, with outline id as a fixed effect (Model:$b$=0.27, p-value=0.01; Outline length: $b$=0.01, p-value=0.30). That is, the length of outlines is not predictive of scores.. We report that evaluating highly subjective tasks may introduce varied ratings among human evaluators despite assigning a large number of evaluators to derive the majority opinion (Krippendorff’s $\alpha$: SFT-QPlanner = 0.22; DPO-QPlanner = 0.23), as observed in other studies (Rottger et al., 2022; Abercrombie et al., 2023).
5.3 Human RAG Evaluation
@@
For long-form response evaluation, we do not automatically measure due to the lack of reliability in long-form text evaluation (Xu et al., 2023). Instead, we recruit another participant to validate the effectiveness of QPlanner on RAG downstream tasks [^8] Details such as recruitment, instructions, and compensation are described in Appendix G.2. 100 RAG responses from the test set are sampled for evaluation, and ten evaluators are assigned for each response. Following insights from Kim et al. (2024b) where the writing format of model responses affects human preferences, we fix the response format with Markdown to compare responses by focusing only on their content. In addition, we prompt LLM999gpt-4-0125-preview is used. to generate responses by strictly relying on given evidence to prevent LLM from arbitrarily responding with its parametric knowledge. We assume web search scenarios for the RAG setup, providing detailed information in Section G.2.1.
@@
Regarding evaluation criteria, we first instruct participants to judge whether generated responses follow requirements of C2 queries or not (Query Satisfaction in Figure 4). We guide them to annotate "Yes" if responses at least partially address topics within C2 queries. For response pairs annotated as "Yes" in both models, participants select their preferred response (Response Preference in Figure 4).
@@
We verify whether subqueries within outlines can help search relevant documents. We compare responses of vanilla RAG with those of DPO-QPlanner using the exactly same prompt. That is, subqueries of DPO-QPlanner only affect the search result. As illustrated in Figure 4(a), we observe that the conventional RAG pipeline does not properly retrieve relevant evidence for answering C2 queries, whose requirements are far more complicated than normal queries. Furthermore, among responses that satisfy the requirements of C2 queries, responses of DPO-QPlanner are mostly preferred.
@@
We further investigate whether better outlines lead to better responses. In this setup, we compare responses of SFT-QPlanner and DPO-QPlanner. The exactly same prompt is used for this comparison, and subqueries within outlines are included in the prompt for composing responses and retrieving documents. Results in Figure 4(b) indicate that further aligning QPlanner with preference can provide preferred outlines, while SFT-QPlanner also shows a similar tendency with DPO-QPlanner.
@@
Figure 4: Human evaluation results.
Since the outcome of both studies is binary (i.e., satisfactory or not), and each human evaluator judges both responses for the same query (within-subjects design), we conduct two McNemar’s tests (McNemar, 1947) to examine whether the differences we find are statistically significant. The contingency tables used for the tests can be found in Table 10. The results confirm that DPO-QPlanner significantly outperforms Vanilla RAG (test-statistics = $60$, p-value < 0.001) and SFT-QPlanner (test-statistics = $207$, p-value = 0.004).
@@ 6 Conclusion
In this work, we suppose complicated user scenarios asking for a constrained range of a specific topic, called coverage-conditioned (C2) query scenarios. To simulate C2 scenarios and a controlled environment for creating and evaluating query outlines (RQ1), we construct QTree, hierarchical sets of subqueries representing diverse perspectives of the original query. Playing a role as boundaries for available outlines, QTree allows systematic comparison of diverse outlines. Subsequently, we train QPlanner which extracts customized outlines from QTree for C2 queries. Regarding our RQ2, our findings based on automatic and human evaluation show that (1) preference-aligned QPlanner can generate better outlines, (2) outlines enable improved document search, and (3) better outlines lead to preferred responses. We believe our work shows the possibility of QTree as a testbed for exploring effective pre-writing strategies to deal with complicated queries.
Limitations
We discuss the current limitations of our work. First, our graphical representation of subquery nodes adheres to canonical tree structures, with each node connected to three child nodes, but it can be adjusted (i.e., composing more or less subqueries) according to tasks or domains. For example, in a complex domain like medical diagnosis, a larger number of subqueries might be necessary to cover various symptoms, possible conditions, diagnostic tests, and treatment options. In contrast, for a straightforward factual query in a domain like mathematics, fewer subqueries might be sufficient to reach a comprehensive answer. Identifying this optimal number still remains an open question and represents a promising direction for future investigation. We believe that our experimental setup serves as an initial testbed for validating these research questions.
It should also be noted that the contents of retrieved documents in our RAG setup can affect the detailed factual consistency of final responses. Although we set the same search configuration among all methodologies, additional fact verification of documents and responses is still needed for practical applications.
While demonstrating significant performance gains both in automatic and human judgements, we find that state-of-the-art LLMs still have difficulty generating long-form responses that handle detailed coverage of C2 queries. This is presumably due to the complexness of C2 queries, and it arises the importance of constructing meticulous benchmarks evaluating long-form responses to complicated queries. This will be another direction of the future work. Lastly, we would like to mention that our five-point scoring schema can be further improved by considering multiple aspects with a fine-grained score rubric.
Ethical Considerations
Since our QTree is generated with benchmarks based on Wikipedia and domain experts, we do not filter sensitive or unsafe contents throughout our studies. For the practical application in the future, deliberate content selection will be required for the safety. In addition, we explicitly share our experimental setup of human studies for transparency in Appendix G.